OpenStudy (anonymous):

Find S8. 3+-6+12-...

OpenStudy (sgadi):

\[a,ar,ar^2 ...\] for our case a=3 r=-2 sum of first n terms is \[\frac{a\left(1-r^{(n+1)}\right)}{1-r}\] so by substituing, we get \[\frac{3(1+128)}{3}=129\]

OpenStudy (dumbcow):

incorrect \[=\frac{a(1-r ^{n}}{1-r}\] \[=\frac{3(1-(-2)^{8})}{3}\] = 1-256 =-255